Building bricks are closely stacked in a pile 7 feet high,36 feet long and 12 feet wide.If the bricks are 2 inches by 4 inches by 9 inches,how many bricks in the pile?
Answer by checkley77(12844)

7*36*12=3,024 cubic feet for the entire pile of bricks.
2*4*9=72 cubic inches for each brick.
There are 12*12*12=1,728 cubic inches in a cubic foot.
1,728/72=24 bricks per cubic foot.
3,024*24=72,576 bricks in the pile.